OpenStudy (anonymous):

How many different ways can the front row seats be filled with 20 students if the classroom has 4 rows of seat?

OpenStudy (lgbasallote):

use combination \[nCr = \frac{n!}{r!(n-r)!}\] \[20C4 = \frac{20!}{4!(20-4)!}\]

OpenStudy (lgbasallote):

do you get it?

OpenStudy (anonymous):

So it would be 4,845?

OpenStudy (lgbasallote):

let me get my scientific calculator

OpenStudy (lgbasallote):

yup.you're right

OpenStudy (anonymous):

Ok, Thanks!

OpenStudy (lgbasallote):

<tips imaginary hat>

OpenStudy (anonymous):

haha :)